How many grams of solute are present in each of these solutions?

a.    37.2 mL of O.471 M HBr




1
Expert's answer
2021-05-06T07:38:29-0400

Volume = 37.2 mL × = 0.0372 L

Moles of solute = 0.471× 0.0372 L = 0.0175 mols

Molar mass of HBr = (1.00 + 80.00) = 81.00 g/mol

Mass of solute = 0.0175 mols × 81.00g/mols = 1.418 g